### v1

# Symmetries of the periodic Fredkin chain

Abstract.
Fredkin chain is a spin-$1/2$ model with interaction of three nearest neighbors. In the case of periodic boundary conditions the ground state is degenerated and can be described in terms of equivalence classes of the Dyck paths. We introduce two operators commuting with the Hamiltonian, which play the roles of lowering and raising operators when acting at the ground states. These operators generate the $B$- or $C$-type Lie algebras, depending on whether the number of sites $N$ is odd or even, respectively, with rank $n=\lceil N/2\rceil$. The third component of total spin operator can be represented as a sum of the Cartan subalgebra elements and some central element. In the $C$-type Lie algebra case (even number of sites) this representation coincides with a similar formula previously conjectured for spin-$1$ operators, in the context of periodic Motzkin chain.
1. Introduction
The Fredkin gate was originally introduced in the context of reversible quantum calculations [FT-82]. It is also known as controlled-SWAP gate and involves three qubits in operation. Relatively recently it was experimentally realized with the help of entangled photons [PHFRP-16].
The Fredkin chain is a model of interacting $s=1/2$ spins generalizing in a certain sense the famous Heisenberg XXX chain [SK-17, SK-18]. In this model, the spin-00 component of two $s=1/2$ spins interacts with the third neighbor (left or right). Under suitably chosen open boundary conditions the ground state of the Hamiltonian can be given as a homogeneous sum of Dyck paths. This state is highly entangled: it shows a logarithmic growth of entanglement entropy with respect to the size of a subsystem. The similar phenomenon is known in the context Motzkin chain, which is model of nearest-neighbor interacting $s=1$ spins, and with the ground state given by a sum of Motzkin paths [BCMNS-12, M-17].
Furthermore, the Fredkin chain appears to be even more related to the Motzkin chain, as it can be viewed as an half-integer spin analogue of the latter. Both models exhibit criticality without a “frustration”: the energy gap vanishes for large system sizes and the unique ground states minimize all individual terms of the Hamiltonians. This property ensures stability of the ground state against inclusion in the Hamiltonian term-dependent interaction parameters. The similarities extend to “colored” versions of these models corresponding to their higher spin generalizations [MS-16, SUZKK-17, ZAK-17]. For these reasons both models are of crucial importance for study of long-range entanglement in critical one-dimensional quantum systems and related topics [DaBT-19, AAZK-19, SVG-23, CBG-24, MGM-24, WZYK-24, VPKGRP-24, CSGV-25].
An interesting question about the Fredkin chain to be raised whether it is an quantum integrable model. The Heisenberg XXX chain is a famous paradigmatic example of such a system [TF-79, S-82, TTF-83, KBI-93, F-96, S-20]. Integrability has been found for a “free” version of the Motzkin chain, with one of the terms describing interaction of spins removed [HSK-23]. Recently, it was observed, albeit purely experimentally for systems of small sizes, that the Motzkin chain with periodic boundary conditions possesses a quite rich symmetry algebra [P-25], that could be a signal of its quantum integrability.
The purpose of the present study is to obtain similar results for the periodic Fredkin chain. Alike for the Motzkin chain, our study here is mainly conjectural, although the central result about properties of the raising and lowering operators is rigorous (see Theorem 2 below). Other results are collected in four conjectures formulated by studying systems of small size, for systems of up to eight sites in the length.
Our main object which we study are operators which play the roles of lowering and raising operators when acting the ground states. These operators decreasing and increasing, respectively, the eigenvalue of the third component of total spin operator, which labels these states. In the route of our study we have also encountered an alternative though less explicit representation for these operators in which they are expressed in terms other components of the total spin operator. This immediately implies that the same operators also commute with the Hamiltonian of the Heisenberg XXX spin-$1/2$ chain. On the other hand, we also conjecture that these operators generate the $B$- or $C$-type Lie algebras, depending on whether the number of sites $N$ is odd or even, respectively, with rank $n=\lceil N/2\rceil$. Furthermore, there exists a central element of the symmetry algebra which can be given as a sum the third component of total spin operator and Cartan subalgebra elements. Surprisingly, we find that in the $C$-type case (even number of sites) this representation coincides with a similar formula previously appeared in the context of periodic Motzkin chain, that hints at the fact that this relation is somewhat representation independent.
The paper is organized as follows. In the next section we review known basic results about the periodic Fredkin chain. In section 3 we introduce the raising and lowering operators and the main result about their properties. In section 4 we study the algebra they generate and show that the third component of total spin operator contains a central element extending this algebra. We conclude in section 5 with a brief discussion of the results and possible proofs of the formulated conjectures.
2. The Hamiltonian, Dyck paths, and the ground states
In this section we recall main ingredients from [SK-17] about the Fredkin chain, with a focus on the periodic boundary conditions.
2.1. The Hamiltonian
We begin with fixing the notation. We denote basis vectors in $\mathbb{C}^{2}$ by
$\left\lvert\uparrow\right\rangle=\begin{pmatrix}1\\ 0\end{pmatrix},\qquad\left\lvert\downarrow\right\rangle=\begin{pmatrix}0\\ 1\end{pmatrix}.$ (2.1)
The basis in $\mathrm{End}(\mathbb{C}^{2})$ is provided by the Pauli matrices
$\sigma^{+}=\begin{pmatrix}0&1\\ 0&0\end{pmatrix},\qquad\sigma^{-}=\begin{pmatrix}0&0\\ 1&0\end{pmatrix},\qquad\sigma^{z}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}.$ (2.2)
We will use also projectors to the spin-up and spin-down states:
$n^{\uparrow}=\frac{1}{2}(1+\sigma^{z}),\qquad n^{\downarrow}=\frac{1}{2}(1-\sigma^{z}).$ (2.3)
The space of states of the $N$-site Fredkin chain is the vector space $(\mathbb{C}^{2})^{\otimes N}$ in which the $j$th factor corresponds to the $j$th site of the spin chain. The basis vectors of $(\mathbb{C}^{2})^{\otimes N}$ are denoted by
$\left\lvert\ell_{1}\ell_{2}\dots\ell_{N}\right\rangle=\left\lvert\ell_{1}\right\rangle\otimes\left\lvert\ell_{2}\right\rangle\otimes\dots\otimes\left\lvert\ell_{N}\right\rangle,\qquad\ell_{1},\ell_{2},\dots,\ell_{N}=\mathrm{\uparrow,\downarrow},$ (2.4)
and by $\sigma_{j}^{\pm,z}$ — operators acting as Pauli matrices in the $j$th copy of $(\mathbb{C}^{2})^{\otimes N}$. For a later reference we need the “total spin operator” for which we choose different normalization of components,
$\mathrm{S}^{\pm}=\sum_{j=1}^{N}\sigma_{j}^{\pm},\qquad\mathrm{S}^{z}=\frac{1}{2}\sum_{j=1}^{N}\sigma_{j}^{z},$ (2.5)
such that the eigenvalues of $\mathrm{S}^{z}$ are $S^{z}=-\frac{N}{2},-\frac{N}{2}+1,\dots,\frac{N}{2}-1,\frac{N}{2}$ (we follow the convention to use upright capital letters for “global” operators and their italic variants for their eigenvalues).
Consider two adjacent copies of $\mathbb{C}^{2}$, say, $j$th and $(j+1)$th. Important objects are the permutation operator $\mathrm{P}_{j,j+1}$ which exchanges these copies, and the projector $\Pi_{j,j+1}$ to the spin-00 state, which is essentially the antisymmetrizer,
$\Pi_{j,j+1}=\frac{1}{2}(1-\mathrm{P}_{j,j+1}).$ (2.6)
The Hamiltonian density of the Fredkin spin chain describes an interaction the spin-00 component of two $s=1/2$ spins with a neighboring third such spin as follows:
$\mathrm{F}_{j,j+1,j+2}=n_{j}^{\uparrow}\,\Pi_{j+1,j+2}+\Pi_{j,j+1}\,n^{\downarrow}_{j+2}.$ (2.7)
The periodic boundary conditions mean that the Hamiltonian is the cyclic sum
$\mathrm{H}=\sum_{j=1}^{N}\mathrm{F}_{j,j+1,j+2},$ (2.8)
where identification of sites by modulus $N$ is assumed.
The operator
$\mathrm{C}=\mathrm{P}_{1,2}\cdots\mathrm{P}_{N-2,N-1}\mathrm{P}_{N-1,N}.$ (2.9)
is the cyclic shift operator $\mathrm{C}$ acting at sites as $i\mapsto i+1$, and it is an integral of motion,
$\left[\mathrm{C},\mathrm{H}\right]=0.$ (2.10)
The operator $\mathrm{S}^{z}$ is also an integral of motion
$\left[\mathrm{S}^{z},\mathrm{H}\right]=0,$ (2.11)
thus eigenvalues of both the cyclic shift operator $\mathrm{C}$ and the third component of total spin operator $\mathrm{S}^{z}$ can be used to label eigenstates of the Hamiltonian.
2.2. Dyck paths
We only summarize the construction developed in [SK-17] to the extent which is sufficient to describe ground states of the Hamiltonian (2.8) below.
Under a Dyck path we will understand a random walk on the square lattice with vertices labeled by $(x,y)$ where $x$ and $y$ are non-negative integers. A Dyck path of length $N$ starts at $(x,y)=(a,0)$ and ends at $(x,y)=(b,N)$, at each step $\Delta x=1$ and $\Delta y=\pm 1$, with the condition that at least once the path hits the $x$-axis, i.e., $\min y=0$ along the path. Note that $a$, $b$, and $N$ are related: there must exist $k\geqslant 0$ such that $N=2k+a+b$. Following [SK-17] we will call all such paths as belonging to the equivalence class $C_{a,b}(N)$. An example is shown in Figure 1.
[…]
Figure 1. A Dyck path belonging to the class $C_{a,b}(N)$, with $a=2$, $b=3$, and $N=9$
The paths in the class $C_{a,b}(N)$ are equivalent up to the recursive procedure where at each iteration a pair of consecutive up and down steps, forming a peak, can be moved anywhere along the path (in [SK-17] these iterations were called Fredkin moves). Thus, one can transform a given Dyck path to some “reference” path which can be chosen, for example, as the path having first $a$ steps down, next $k$ pairs of peaks, and, finally, $b$ steps up.
Apparently, all possible Dyck paths of length $N$ are in a one-to-one correspondence with the basis vectors of $(\mathbb{C}^{2})^{\otimes N}$, the state space of the Fredkin chain. The correspondence is based on a simple observation that for each letter $\ell_{j}$ in (2.4) one can associate step up or step down according to whether $\ell_{j}=\uparrow$ or $\ell_{j}=\downarrow$. Thus, a string (or, a “word”) $\ell_{1}\dots\ell_{N}$ uniquely defines the Dyck path, and this can be used for their notation.
Division of Dyck paths on equivalence classes introduces certain order among the basis vectors. In particular, vectors associated with the class $C_{a,b}(N)$ are eigenvectors of $\mathrm{S}^{z}$ with the eigenvalue $S^{z}=\frac{1}{2}(b-a)$. Furthermore, one can introduce states $\left\lvert C_{a,b}(N)\right\rangle$ as homogeneous sums over Dyck paths belonging to $C_{a,b}(N)$, namely,
$\left\lvert C_{a,b}(N)\right\rangle=\sum_{\ell_{1}\ell_{2}\dots\ell_{N}\in C_{a,b}(N)}\left\lvert\ell_{1}\ell_{2}\dots\ell_{N}\right\rangle.$ (2.12)
For example, in the case $N=3$ the allowed Dyck paths, see Figure 2, yield
$\left\lvert C_{0,3}(3)\right\rangle=\left\lvert\uparrow\uparrow\uparrow\right\rangle,\qquad\left\lvert C_{0,1}(3)\right\rangle=\left\lvert\uparrow\downarrow\uparrow\right\rangle+\left\lvert\uparrow\uparrow\downarrow\right\rangle,\qquad\left\lvert C_{1,2}(3)\right\rangle=\left\lvert\downarrow\uparrow\uparrow\right\rangle,$ (2.13)
and there are also the states $\left\lvert C_{3,0}(3)\right\rangle$, $\left\lvert C_{1,0}(3)\right\rangle$, and $\left\lvert C_{2,1}(3)\right\rangle$ obtained by reversal of arrows of these states (mirror transformation of paths), respectively.
$C_{0,3}(3)$$C_{0,1}(3)$$C_{1,2}(3)$
Figure 2. Equivalence classes of Dyck paths in the case $N=3$; the remaining can be obtained by a mirror transformation
The Dyck paths for $N$ even belonging to the class $C_{0,0}(N)$ correspond to the Dyck paths in their standard definition, i.e., these are paths that start at $(0,0)$ and ends at $(0,N)$ with the condition that they are allowed to pass below the $x$-axis. The corresponding state $\left\lvert C_{0,0}(N)\right\rangle$ is a unique ground state of the open Fredkin chain introduced in [SK-17, SK-18].
2.3. Ground states
In the case of periodic boundary conditions, the ground states corresponding to zero eigenvalue of the Hamiltonian can be constructed as superposition of states $\left\lvert C_{a,b}(N)\right\rangle$ with $a$ and $b$ related to an eigenvalue of the operator $\mathrm{S}^{z}$. A key observation leading to expressions for the ground states is an extension of the Fredkin moves by periodicity, see [SK-17, SK-18], for details. To emphasize that this an established result, we formulate it here as a theorem.
Theorem 1 (Salberger, Korepin [SK-17, SK-18]).
For each value of $S^{z}\in\{-\frac{N}{2},-\frac{N}{2}+1,\dots,\frac{N}{2}-1,-\frac{N}{2}\}$ there exists one and only one ground state $\left\lvert v_{S^{z}}\right\rangle$ of the periodic Fredkin chain Hamiltonian (2.8) with zero eigenvalue, which correspond to the eigenvalue $C=1$ of the cyclic shift operator, of the form
$\left\lvert v_{S^{z}}\right\rangle=\sum_{\begin{subarray}{c}a,b\geqslant 0\\ b-a=2S^{z}\end{subarray}}\left\lvert C_{a,b}(N)\right\rangle.$ (2.14)
In the case of $N$ even, $N=2n$, and $S^{z}=0$ there exists one more ground state which correspond to the eigenvalue $C=-1$, of the form
$\left\lvert v_{0}^{-}\right\rangle=\sum_{k=0}^{n}(-1)^{k}\left\lvert C_{k,k}(2n)\right\rangle.$ (2.15)
Thus [^1] In [SK-17, SK-18] the degeneracies for odd and even $N$ were misprinted as $N$ and $N+1$, respectively., in the case of $N$ odd the degeneracy of the ground state is $N+1$ and in the case of $N$ even is $N+2$.
$C_{0,0}(4)$$C_{2,2}(4)$
$C_{1,1}(4)$
Figure 3. Three classes $C_{a,a}(N)$ of the Dyck paths for $N=4$; the first line (even $a$) gives rise to the state $\left\lvert v_{0}^{\text{even}}\right\rangle$, and the second line (odd $a$) to the state $\left\lvert v_{0}^{\text{odd}}\right\rangle$
Appearance of two ground states for $N$ even and $S^{z}=0$ comes from the fact that the considerations involving the periodic Fredkin moves lead to existence of two eigenstates, with even and odd values of $a$, for the classes $C_{a,a}(N)$. An example of $N=4$ is shown in Figure 3. These states are
$\left\lvert v_{0}^{\text{even}}\right\rangle=\sum_{k=0}^{n}\left\lvert C_{2k,2k}(2n)\right\rangle,\qquad\left\lvert v_{0}^{\text{odd}}\right\rangle=\sum_{k=0}^{n-1}\left\lvert C_{2k+1,2k+1}(2n)\right\rangle,$ (2.16)
and they can be mapped to each other by the cyclic shift operator,
$\mathrm{C}\left\lvert v_{0}^{\text{even}}\right\rangle=\left\lvert v_{0}^{\text{odd}}\right\rangle,\qquad\mathrm{C}\left\lvert v_{0}^{\text{odd}}\right\rangle=\left\lvert v_{0}^{\text{even}}\right\rangle.$ (2.17)
Their linear combinations $\left\lvert v_{0}^{\pm}\right\rangle=\left\lvert v_{0}^{\text{even}}\right\rangle\pm\left\lvert v_{0}^{\text{odd}}\right\rangle$ are eigenstates of the cyclic shift operator with the eigenvalues $C=\pm 1$, thus leading to the states $\left\lvert v_{0}\right\rangle\equiv\left\lvert v_{0}^{+}\right\rangle$ and $\left\lvert v_{0}^{-}\right\rangle$ mentioned above.
3. Lowering and raising operators
Now we turn to the results. We have obtained them mainly by a direct study of the model for up to $N=8$. The central result, about explicit expressions for raising and lowering operators, is given here as a theorem. The remaining results (in this and the next section) are summarized in four conjectures.
3.1. Symmetry generators
In [SK-17, SK-18] it was indicated that the states $\left\lvert v_{S^{z}}\right\rangle$ are nothing but the ground states of the spin-$1/2$ ferromagnetic Heisenberg XXX chain (see, e.g., [F-96]). As such, they can be obtained from each other by acting with the operators $\mathrm{S}^{+}$ and $\mathrm{S}^{-}$, which increase and decrease, respectively, the value of $S^{z}$. However, these operators has no similar role in the context of the Fredkin chain since they do not commute with the Hamiltonian (2.8).
One of our main results here is that there exists another pair of operators which play the similar role of raising and lowering operators when acting at the states $\left\lvert v_{S^{z}}\right\rangle$, but commuting with the Hamiltonian (2.8).
Theorem 2.
The operators
$\Sigma^{\pm}=\sum_{\begin{subarray}{c}r_{1},\dots,r_{N}\in\{-1,0,1\}\\ r_{1}+\dots+r_{N}=\pm 1\end{subarray}}\sigma_{1}^{r_{1}}\cdots\sigma_{N}^{r_{N}},$ (3.1)
where $\sigma_{i}^{0}=1$ and $\sigma_{i}^{\pm 1}=\sigma_{i}^{\pm}$, or, equivalently,
$\Sigma^{\pm}=\operatorname*{\mathrm{res}}_{\lambda=0}\,\prod_{i=1}^{N}\Bigl{(}\lambda^{-1}\sigma_{i}^{\pm}+1+\lambda\sigma_{i}^{\mp}\Bigr{)},$ (3.2)
when acting at the ground states $\left\lvert v_{S^{z}}\right\rangle$ play the role of raising and lowering operators, respectively,
$\Sigma^{\pm}\left\lvert v_{S^{z}}\right\rangle=\begin{cases}c_{\pm}(S^{z})\left\lvert v_{S^{z}\pm 1}\right\rangle,&S^{z}\neq\pm N/2,\\ 0,&S^{z}=\pm N/2,\end{cases}$ (3.3)
where $c_{\pm}(S^{z})\neq 0$ are some constants, and commute with the Hamiltonian,
$\left[\Sigma^{\pm},\mathrm{H}\right]=0.$ (3.4)
We give a proof below at the end of this section. Surprisingly, the assertion of the theorem about commutativity with the Hamiltonian can be seen as a corollary of the technically stronger statement (see Proposition 1 below).
The operators $\Sigma^{\pm}$ annihilate the state $\left\lvert v_{0}^{-}\right\rangle$ specific for $N$ even,
$\Sigma^{\pm}\left\lvert v_{0}^{-}\right\rangle=0.$ (3.5)
This property can be seen as a particular case of the following more general observation.
Conjecture 1.
The operators $\Sigma^{\pm}$ annihilate all non-cyclic invariant ($C\neq 1$) eigenstates of the Hamiltonian.
3.2. A relation to the total spin operator
As it clear from (3.1), the matrices representing operators $\Sigma^{\pm}$ are build of 00’s and $1$’s; these matrices are nilpotent, $(\Sigma^{\pm})^{N+1}=0$, though not strictly upper or lower triangle ones. The operators $\Sigma^{\pm}$ are non-local, contrary to the operators $\mathrm{S}^{\pm}$, and at a first glance, they are unrelated. However, it turn out that $\Sigma^{\pm}$’s can be expressed in terms of $\mathrm{S}^{\pm}$’s.
Conjecture 2.
For the operators $\Sigma^{\pm}$ there exists the representation
$\Sigma^{\pm}=\sum_{k=1}^{\lceil N/2\rceil}\gamma_{k}\left(\operatorname*{\mathrm{aad}}\mathrm{S}^{\pm}\operatorname*{\mathrm{aad}}\mathrm{S}^{\mp}\right)^{k-1}\mathrm{S}^{\pm},$ (3.6)
where $\gamma_{k}$ are some coefficients and $\operatorname*{\mathrm{aad}}$ denotes the anti-adjoint action, $\operatorname*{\mathrm{aad}}a\equiv\{a,\cdot\,{}\}$, where braces stand for the anti-commutator.
For example, for $N=3$,
$\Sigma^{\pm}=-\frac{1}{4}\mathrm{S}^{\pm}+\frac{1}{8}\{\mathrm{S}^{\pm},\{\mathrm{S}^{\mp},\mathrm{S}^{\pm}\}\}.$ (3.7)
Coefficients $\gamma_{k}$ for the cases of $N$ up to $N=8$ are
$\begin{gathered}\gamma_{1}=-\frac{1}{4},\quad\gamma_{2}=\frac{1}{8},\qquad N=3,\\ \gamma_{1}=-\frac{3}{4},\quad\gamma_{2}=\frac{1}{8},\qquad N=4,\\ \gamma_{1}=\frac{1}{8},\quad\gamma_{2}=-\frac{7}{96},\quad\gamma_{3}=\frac{1}{192},\qquad N=5,\\ \gamma_{1}=\frac{5}{8},\quad\gamma_{2}=-\frac{13}{96},\quad\gamma_{3}=\frac{1}{192},\qquad N=6,\\ \gamma_{1}=-\frac{5}{64},\quad\gamma_{2}=\frac{37}{768},\quad\gamma_{3}=-\frac{11}{2304},\quad\gamma_{4}=\frac{1}{9216},\qquad N=7,\\ \gamma_{1}=-\frac{35}{64},\quad\gamma_{2}=\frac{101}{768},\quad\gamma_{3}=-\frac{17}{2304},\quad\gamma_{4}=\frac{1}{9216},\qquad N=8.\end{gathered}$ (3.8)
The values of $\gamma_{1}$ at $N=7$ and $N=8$ breaks an easy guess that they should differ be $1/2$, as it happens at $N=3$ and $N=4$, and at $N=5$ and $N=6$.
3.3. Proof of Theorem 2
To show that (3.3) holds, we start with rewriting one of the results of Theorem 1, namely, (2.14) in a form suitable for application of formula (3.1) for the operators $\Sigma^{\pm}$.
Lemma 1.
The cyclic-invariant ($C=1$) ground states $\left\lvert v_{S^{z}}\right\rangle$ can be written as follows:
$\left\lvert v_{N/2-m}\right\rangle=\sum_{1\leqslant k_{1}<k_{2}<\dots<k_{m}\leqslant N}\sigma^{-}_{k_{1}}\sigma^{-}_{k_{2}}\cdots\sigma^{-}_{k_{m}}\left\lvert\Uparrow\right\rangle.$ (3.9)
Equivalently,
$\left\lvert v_{N/2-m}\right\rangle=\operatorname*{\mathrm{res}}_{\lambda=0}\lambda^{m-1}\prod_{i=1}^{N}\left(\lambda^{-1}\sigma^{-}_{i}+1\right)\left\lvert\Uparrow\right\rangle$ (3.10)
or
$\left\lvert v_{N/2-m}\right\rangle=\operatorname*{\mathrm{res}}_{\lambda=0}\lambda^{-m-1}\prod_{i=1}^{N}\left(\lambda\sigma^{-}_{i}+1\right)\left\lvert\Uparrow\right\rangle.$ (3.11)
[…]
Formula (3.9) is simply (2.14) written in terms of the Pauli spin operators; (3.10) and (3.11) are analogues or (3.2).
Now, one can prove the result of action of the operators $\Sigma^{\pm}$ at the vectors $\left\lvert v_{S^{z}}\right\rangle$. Using the residue formulas (3.2) and (3.10), we find
$\Sigma^{\pm}\left\lvert v_{N/2-m}\right\rangle$ $=\operatorname*{\mathrm{res}}_{\lambda=0}\operatorname*{\mathrm{res}}_{\mu=0}\lambda^{m-1}\prod_{i=1}^{N}\left(\mu^{-1}\sigma_{i}^{\pm}+1+\mu\sigma_{i}^{\mp}\right)\left(\lambda^{-1}\sigma_{i}^{-}+1\right)\left\lvert\Uparrow\right\rangle$ (3.12)
$=\operatorname*{\mathrm{res}}_{\lambda=0}\operatorname*{\mathrm{res}}_{\mu=0}\lambda^{m-1}\prod_{i=1}^{N}\left((\mu^{\pm 1}+\lambda^{-1})\sigma_{i}^{-}+1+\mu^{\mp 1}\lambda^{-1}\sigma_{i}^{+}\sigma_{i}^{-}\right)\left\lvert\Uparrow\right\rangle$ (3.13)
$=\operatorname*{\mathrm{res}}_{\lambda=0}\operatorname*{\mathrm{res}}_{\mu=0}\lambda^{m-1}\left(1+\mu^{\mp 1}\lambda^{-1}\right)^{N}\prod_{i=1}^{N}\left(\mu^{\pm 1}\sigma_{i}^{-}+1\right)\left\lvert\Uparrow\right\rangle$ (3.14)
$=\binom{N}{m}\operatorname*{\mathrm{res}}_{\mu=0}\mu^{\pm m}\prod_{i=1}^{N}\left(\mu^{\pm 1}\sigma_{i}^{-}+1\right)\left\lvert\Uparrow\right\rangle,$ (3.15)
and (3.3) follows, after taking into account (3.10) (for $\Sigma^{-}$) or (3.11) (for $\Sigma^{+}$).
Let us now consider commutativity of the operators $\Sigma^{\pm}$ with the Hamiltonian, relation (3.4). To this end it is useful to consider more general operators:
$\Sigma^{Q}_{i_{1},\dots,i_{n}}=\sum_{\begin{subarray}{c}r_{1},\dots,r_{n}\in\{-1,0,1\}\\ r_{1}+\dots+r_{n}=Q\end{subarray}}\sigma_{i_{1}}^{r_{1}}\cdots\sigma_{i_{n}}^{r_{n}},\qquad Q=0,\pm 1,\dots,\pm n.$ (3.16)
Here, as above, $\sigma_{i}^{0}\equiv 1$, $\sigma_{i}^{\pm 1}\equiv\sigma_{i}^{\pm}$, and the subscripts $i_{1},\dots,i_{n}$, indicate copies of $\mathbb{C}^{2}$ where the Pauli spin operators act, $1\leqslant i_{1}<\dots<i_{n}\leqslant N$. The superscript $Q$ can be regarded as a “charge” of the operator. Note that with this notation $\Sigma^{\pm}\equiv\Sigma^{\pm 1}_{1,\dots,N}$, and we will follow the convention to drop the subscripts when $n=N$, i.e., writing $\Sigma^{Q}\equiv\Sigma^{Q}_{1,\dots,N}$. The following observation is crucial.
Proposition 1.
The Hamiltonian density of the Fredkin chain (2.7) satisfies
$[\Sigma^{Q},\mathrm{F}_{i,i+1,i+2}]=0,\qquad Q=0,\pm 1,\dots,\pm N.$ (3.17)
Proof.
Because of the cyclic invariance of the operators $\Sigma^{Q}$, to prove (3.17) it is sufficient to consider the case $i=1$, i.e., to show that it holds for the operator $\mathrm{F}_{123}$. The operators $\Sigma^{Q}$ admit the decomposition
$\Sigma^{Q}=\sum_{\tilde{Q}=-3}^{3}\Sigma_{123}^{\tilde{Q}}\Sigma_{4,\dots,N}^{Q-\tilde{Q}}.$ (3.18)
Note that this decomposition simply follows from the core definition (3.16) and it can also be derived formally by writing a residue formula by analogue with (3.2), and next evaluating the residues after the summation. The decomposition (3.18) implies that (3.17) will follow if $[\Sigma_{123}^{Q},\mathrm{F}_{123}]=0$ for $Q=0,\pm 1,\pm 2,\pm 3$. Explicitly, the operators appearing here are
$\Sigma^{0}_{123}$ $=1+\sigma_{1}^{+}\sigma_{2}^{-}+\sigma_{1}^{-}\sigma_{2}^{+}+\sigma_{1}^{+}\sigma_{3}^{-}+\sigma_{1}^{-}\sigma_{3}^{+}+\sigma_{2}^{+}\sigma_{3}^{-}+\sigma_{2}^{-}\sigma_{3}^{+},$ (3.19)
$\Sigma^{\pm 1}_{123}$ $=\sigma_{1}^{\pm}+\sigma_{2}^{\pm}+\sigma_{3}^{\pm}+\sigma_{1}^{\pm}\sigma_{2}^{\pm}\sigma_{3}^{\mp}+\sigma_{1}^{\pm}\sigma_{2}^{\mp}\sigma_{3}^{\pm}+\sigma_{1}^{\mp}\sigma_{2}^{\pm}\sigma_{3}^{\pm},$ (3.20)
$\Sigma^{\pm 2}_{123}$ $=\sigma_{1}^{\pm}\sigma_{2}^{\pm}+\sigma_{1}^{\pm}\sigma_{3}^{\pm}+\sigma_{2}^{\pm}\sigma_{3}^{\pm},$ (3.21)
$\Sigma^{\pm 3}_{123}$ $=\sigma_{1}^{\pm}\sigma_{2}^{\pm}\sigma_{3}^{\pm}.$ (3.22)
A direct check shows that $\mathrm{F}_{123}$ indeed commutes with all of them. ∎
Thus, the commutativity property (3.4) follows simply from a particular case of (3.17) at $Q=\pm 1$. This completes the proof of Theorem 2.
4. Symmetry algebra
Here, we focus on further discussion of properties of the operators $\Sigma^{\pm}$, such as which algebra they span and what is a role of the operator $\mathrm{S}^{z}$ in this construction. We start with comments about one more operator which exists in the case of $N$ even and next proceed to the results.
4.1. Preliminaries
In the case of $N$ even, one can introduce one more operator, $\Xi$, which when acting at the ground states maps the states $\left\lvert v_{0}\right\rangle\equiv\left\lvert v_{0}^{+}\right\rangle$ and $\left\lvert v_{0}^{-}\right\rangle$ to each other,
$\Xi\left\lvert v_{0}^{\pm}\right\rangle=\left\lvert v_{0}^{\mp}\right\rangle,$ (4.1)
and annihilates the states $\left\lvert v_{S^{z}}\right\rangle$ with $S^{z}\neq 0$,
$\Xi\left\lvert v_{S^{z}}\right\rangle=0,\qquad S^{z}=\pm 1,\dots,\pm N/2.$ (4.2)
Inspecting the structure of the states (2.15), it is not difficult to figure out that such an operator can be written in the form
$\Xi=\sum_{k=0}^{N/2}(-1)^{k}\sum_{\ell_{1}\ell_{2}\dots\ell_{N}\in C_{k,k}(N)}n_{1}^{\ell_{1}}n_{2}^{\ell_{2}}\dots n_{N}^{\ell_{N}},$ (4.3)
where $n_{j}^{\ell_{j}}$, $\ell_{j}=\uparrow,\downarrow$, are the spin-up and spin-down projection operators, and the inner summation is performed over the Dyck paths from the class $C_{k,k}(N)$.
As it is clear from (4.3), the operator $\Xi$ is represented by a diagonal matrix with 00’s, $1$’s and $-1$’s in its entries, with 00’s responsible for cutting out all sectors with $S^{z}=\pm 1,\dots,\pm N$, whereas in the sector with $S^{z}=0$ it has only $1$’s and $-1$’s in an equal number. In other words, $\Xi$ acts nontrivially only in the $S^{z}=0$ sector of the whole Hilbert space of the Hamiltonian. It commutes with the Hamiltonian and anti-commutes with the cyclic shift operator:
$[\Xi,\mathrm{H}]=0,\qquad\mathrm{C}\Xi\mathrm{C}^{-1}=-\Xi.$ (4.4)
The first relation may require a proof though we find it satisfied in all checks; the second one is obvious from (4.3). These relations mean that $\Xi$ maps every state in the $S^{z}=0$ sector to another state in this sector with the same eigenvalue of $\mathrm{H}$, but alternated eigenvalue of the cyclic shift operator, $C\mapsto-C$. Existence of such an unusual operator may explain a clear visible (in numerical checks) double degeneracy of the whole spectrum of the Hamiltonian. It complements, for the case of $S^{z}=0$, doubling of states in the $S^{z}\neq 0$ sectors following from the spin-reversal symmetry.
Nevertheless, we exclude $\Xi$ from our considerations below, since in treating the symmetry algebra of the model only operators which commute with both integrals of motion, $\mathrm{H}$ and $\mathrm{C}$, are of interest.
4.2. The Lie algebra of raising and lowering operators
Now we focus on establishing an algebra the operators $\Sigma^{\pm}$ generate. We denote this algebra by $\mathfrak{g}$. Our findings can be summarized as follows.
Conjecture 3.
The rank of $\mathfrak{g}$ is $n=\lceil N/2\rceil$ and
$\mathfrak{g}=\begin{cases}B_{n}=\mathfrak{so}_{2n+1},&N=2n-1,\\ C_{n}=\mathfrak{sp}_{2n},&N=2n.\end{cases}$ (4.5)
To clarify the statement, it is useful to recall (see, e.g., [K-02, IR-18]) that a complex semisimple Lie algebra has a standard definition in terms of Chevalley generators $e_{i},f_{i},h_{i}$, $i=1,\dots,n$, where $n$ is the rank of the algebra, which satisfy the Serre relations:
$\begin{gathered}=0,\qquad[e_{i},f_{i}]=h_{i},\qquad[e_{i},f_{j}]=0,\quad i\neq j,\\ [h_{i},e_{j}]=A_{ij}e_{j},\qquad[h_{i},f_{j}]=-A_{ij}f_{j},\\ (\operatorname*{\mathrm{ad}}e_{i})^{1-A_{ij}}e_{j}=(\operatorname*{\mathrm{ad}}f_{i})^{1-A_{ij}}f_{j}=0,\quad i\neq j.\end{gathered}$ (4.6)
Here, $\operatorname*{\mathrm{ad}}a\equiv[a,\cdot\,{}]$, and the numbers $A_{ij}$ are entries of the Cartan matrix. One can identify the algebra by computing the Cartan matrix that involves construction of the Chevalley generators.
As in [P-25], to formulate the above conjecture we have performed the following calculation. In addition to $\Sigma^{\pm}$, we define $\Sigma^{z}=[\Sigma^{+},\Sigma^{-}]$ and construct triples $\Lambda^{\pm,z}$, $\Phi^{\pm,z}$, etc, introducing them recursively, first $\Lambda^{\pm}=\pm\left[\Sigma^{z},\Sigma^{\pm}\right]$, $\Lambda^{z}=\left[\Lambda^{+},\Lambda^{-}\right]$, next $\Phi^{\pm}=\pm\left[\Lambda^{z},\Lambda^{\pm}\right]$, $\Phi^{z}=\left[\Phi^{+},\Phi^{-}\right]$, and so on. An important observation which can be easily done at this stage is that the operators $\Sigma^{z},\Lambda^{z},\Phi^{z},\dots$ form an abelian algebra, and therefore they can considered as linear combinations of the Cartan subalgebra elements $h_{i}$. Hence, the rank of the algebra can be identified by how many operators $\Sigma^{z},\Lambda^{z},\Phi^{z},\dots$ are linearly independent.
As the rank is found, the next step is to search elements $e_{i}$ and $f_{i}$ in the form
$e_{i}=\rho_{i}\left(\Sigma^{+}+a_{i}\Lambda^{+}+b_{i}\Phi^{+}+\dots\right),\qquad f_{i}=\rho_{i}\left(\Sigma^{-}+a_{i}\Lambda^{-}+b_{i}\Phi^{-}+\dots\right),$ (4.7)
where the number of terms is equal to the rank. Solving relations $[e_{i},f_{j}]=0$, $i\neq j$, for the coefficients $a_{i},b_{i},\dots$ one finds a unique solution up to a permutation of triples of Chevalley generators. The solution is fixed by finding $\rho_{i}$ from the relations $[h_{i},e_{i}]=2e_{i}$, that corresponds to the canonical normalization of the Cartan matrix with $A_{ii}=2$.
At the last step, the off-diagonal entries of the Cartan matrix are computed according to the relations $[h_{i},e_{j}]=A_{ij}e_{j}$. A particular order among the triples of Chevalley generators should be fixed, such that the Cartan matrix can be identified according to the well-known classification by Dynkin diagrams. In our problem we meet Cartan matrices for the Lie algebras of $B$- and $C$-types. Recall that the off-diagonal nonzero entries of the Cartan matrix, in the case of $B$-type algebra are
$A_{i+1,i}=-1,\qquad A_{i,i+1}=-1\quad(i\neq n-1),\qquad A_{N-1,N}=-2,$ (4.8)
and in the case of $C$-type algebra are
$A_{i+1,i}=-1\quad(i\neq n-1),\qquad A_{N,N-1}=-2,\qquad A_{i,i+1}=-1.$ (4.9)
As a self-consistency check, the relations $(\operatorname*{\mathrm{ad}}e_{i})^{1-A_{ij}}e_{j}=0$, $i\neq j$, are verified.
4.3. A central element
Noting that $[\mathrm{S}^{z},\Sigma^{\pm}]=\pm\Sigma^{\pm}$ and one can expect the operator $\mathrm{S}^{z}$ can be expressed terms of the Cartan subalgebra elements. It turns out that such a construction involves one more operator which appears to be a central element with respect to the generators of $\mathfrak{g}$. This can be formulated as follows.
Conjecture 4.
There exists a nontrivial operator $p$ which is a central element with the respect to the algebra $\mathfrak{g}$, and it has the form
$p=\mathrm{S}^{z}-\sum_{i=1}^{\lceil N/2\rceil}\alpha_{i}h_{i},$ (4.10)
where $\mathrm{S}^{z}$ is the third component of total spin operator, $\alpha_{i}$ are some coefficients, and $h_{i}$ are the Cartan subalgebra elements.
We call $p$ central element since it commutes with the operators $\Sigma^{\pm}$ and hence with all the Chevalley generators, $[p,e_{i}]=[p,f_{i}]=[p,h_{i}]$.
Formula (4.10) can be viewed as expression for the operator $\mathrm{S}^{z}$ in terms of elements of $\mathfrak{g}$ and the operator $p$. Examples up to $N=8$ read
$\begin{gathered}\alpha_{1}=2,\quad\alpha_{2}=\frac{3}{2},\qquad N=3,4,\\ \alpha_{1}=\frac{5}{2},\quad\alpha_{2}=4,\quad\alpha_{3}=\frac{9}{2},\qquad N=5,\\ \alpha_{1}=3,\quad\alpha_{2}=5,\quad\alpha_{3}=3,\qquad N=6,\\ \alpha_{1}=\frac{7}{2},\quad\alpha_{2}=6,\quad\alpha_{3}=\frac{15}{2},\quad\alpha_{4}=8,\qquad N=7,\\ \alpha_{1}=4,\quad\alpha_{2}=7,\quad\alpha_{3}=9,\quad\alpha_{4}=5,\qquad N=8.\end{gathered}$ (4.11)
Coincidence of the coefficients in the cases of $N=3$ and $N=4$ can be explained by appearance of the same algebra ($B_{2}=C_{2}$) in these two cases. Furthermore, an interesting conclusion which can be made by inspecting these examples, is that in the case of $C$-type Lie algebra ($N$ even) formula (4.10) exactly coincides with that obtained in [P-25] in the context of the periodic Motzkin chain. This suggests, that the relation (4.10) is probably independent of the choice of the initial representation of the $\mathfrak{sl}_{2}$ algebra of local spin operators.
Inspecting (4.11) one can make some further conjectures. For example, one can guess that $\alpha_{1}=N/2$ for $N>3$, and $\alpha_{2}=N-1$ for $N>4$. However, to get more insight on these numbers, a more output is needed for further values of $N$. This is not a quite simple task since it requires an explicit construction of the Chevalley generators in each particular case, to establish the corresponding relation (4.10).
5. Conclusion
In this paper, we have considered the Fredkin spin chain with periodic boundary conditions basing on the result of Salberger and Korepin [SK-17, SK-18] about explicit form of the ground states (Theorem 1).
We have showed that there exist raising and lowering operators which maps these states to each other and which commute with the Hamiltonian (Theorem 2). We have also provided several conjectures about other properties of these operators. Conjecture 1 says that these operators actually act non-trivially only in the subspace of cyclic invariant states. Conjecture 2 shows that these operators can be (somewhat unexpectedly) be expressed in terms of the non-diagonal components of total spin operator. Conjecture 3 states that they span in $B$- or $C$-type Lie algebra, according to the parity of number of sites of the chain. Conjecture 4 asserts that the symmetry algebra of the model is actually wider and extended by a central element contained in the third component of total spin operator.
In our previous study devoted to the Motzkin chain [P-25] the result analogous to Theorem 2 was formulated as a conjecture. In fact, the proof we give here for Theorem 2 can be directly generalized to the spin-$1$ case of Motzkin chain. This is intimately related to fact that the lowering an raising operators have essentially the same structure for both models.
Conjecture 1 can also be formulated in the context of Motzkin chain. It is not however clear how it can proven in either case. Concerning Conjecture 2, it seems plausible that it can be proven be suitably applying the technique exposed in [ZF-11]. Conjectures 3 and 4 are direct analogues of the corresponding conjectures made in the context of Motzkin chain in [P-25]. We have observed that the central element for the Fredkin chain with even number of sites appears to be related to the third component of total spin operator by exactly the same relation as it was found in the context of the Motzkin chain. In both cases the symmetry algebra is the $C$-type Lie algebra of the same rank. It would be interesting to understand group theoretical backgrounds of these results. In addition, the coefficients $\gamma_{i}$ and $\alpha_{i}$ appearing in Conjecture 2 and Conjecture 4, respectively, certainly have some combinatorial interpretations that can also be a subject of further study.
Acknowledgments
The author is indebted for stimulating discussions to Nikolai M. Bogoliubov, Vladimir E. Korepin, and Vitaly O. Tarasov.